OpenStudy (anonymous):

What is the square root of 160 with -4 outside?

OpenStudy (anonymous):

\[-4\sqrt{160} ?\]

OpenStudy (anonymous):

\[-4\sqrt{4*4*5*2}\]

OpenStudy (anonymous):

\[-4(4)\sqrt{10}\]

OpenStudy (anonymous):

Yes,thanks